### v1

# Examining Kempe equivalence via commutative algebra

Abstract.
Kempe equivalence is a classical and important notion on vertex coloring in graph theory. In the present paper, we introduce several ideals associated with graphs and provide a method to determine whether two $k$-colorings are Kempe equivalent via commutative algebra. Moreover, we give a way to compute all $k$-colorings of a graph up to Kempe equivalence by virtue of the algebraic technique on Gröbner bases. As a consequence, the number of $k$-Kempe classes can be computed by using Hilbert functions. Finally, we introduce several algebraic algorithms related to Kempe equivalence.
Key words and phrases:
vertex coloring, Kempe switching, Kempe equivalence, Gröbner basis, Hilbert function
2020 Mathematics Subject Classification:
05C15, 13P10, 13F65
1. Introduction
A $k$-coloring $f$ of a graph $G$ on the vertex set $[d]:=\{1,2,\ldots,d\}$ is a map from $[d]$ to $[k]$ such that $f(i)\neq f(j)$ for all $\{i,j\}\in E(G)$. The smallest integer $\chi(G)$ such that $G$ has a $\chi(G)$-coloring is called the chromatic number of $G$. Given a $k$-coloring $f$ of $G$, and integers $1\leq i<j\leq k$, let $H$ be a connected component of the induced subgraph of $G$ on the vertex set $f^{-1}(i)\cup f^{-1}(j)$. Then we can obtain a new $k$-coloring $g$ of $G$ by setting
$g(x)=\begin{cases}f(x)&x\notin H,\\ i&x\in H\mbox{ and }f(x)=j,\\ j&x\in H\mbox{ and }f(x)=i.\\ \end{cases}$
We say that $g$ is obtained from $f$ by a Kempe switching. Two $k$-colorings $f$ and $g$ of $G$ are called Kempe equivalent, denoted by $f\sim_{k}g$, if there exists a sequence $f_{0},f_{1},\ldots,f_{s}$ of $k$-colorings of $G$ such that $f_{0}=f$, $f_{s}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching. Let denote $\mathcal{C}_{k}(G)$ the set of all $k$-colorings of $G$. Then $\sim_{k}$ is an equivalence relation on $\mathcal{C}_{k}(G)$. The equivalence classes of $\mathcal{C}_{k}(G)$ by $\sim_{k}$ are called the $k$-Kempe classes. We denote ${\rm kc}(G,k)$ the quotient set $\mathcal{C}_{k}(G)/\sim_{k}$ and denote ${\rm Kc}(G,k)$ the number of $k$-Kempe classes of $G$, namely ${\rm Kc}(G,k)=|{\rm kc}(G,k)|$. See [5] for an overview of Kempe equivalence.
Given a graph $G$, let $S(G)$ be the set of all stable sets of $G$. The stable set ideal $I_{G}$ of $G$ is a toric ideal arising from $S(G)$ of a polynomial ring $R[G]:={\mathbb{K}}[x_{S}\mid S\in S(G)]$ over a field ${\mathbb{K}}$. In [7], the authors showed that $I_{G}$ is generated by binomials ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ associated with $k$-colorings $f$ and $g$ of a replication graph of an induced subgraph of $G$, and found a relationship between Kempe equivalence on $G$ and an algebraic property of $I_{G}$. In particular, by using the proof of [7, Theorem 1.3], we can examine if two $k$-colorings of $G$ are Kempe equivalent by using $I_{G}$. However, $I_{G}$ has too much information for this purpose. In the present paper, we introduce a simpler ideal $J_{G}$, which is generated by binomials ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ associated with $2$-colorings $f$ and $g$ of an induced subgraph of $G$, to determine Kempe equivalence on $G$. We call $J_{G}$ the $2$-coloring ideal of $G$. Then our first main result is the following:
Theorem 1.1.
Let $G$ be a graph on $[d]$ and let $f,g$ be $k$-colorings of $G$. Then $f\sim_{k}g$ if and only if ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$.
Next, we compute all $k$-colorings of a graph $G$ up to Kempe equivalence by virtue of the algebraic technique on Gröbner bases. Namely, a complete representative system for ${\rm kc}(G,k)$ is given. For this, we introduce another ideal $K_{G}$ defined by
$K_{G}:=J_{G}+M_{G},$
where
$M_{G}:=\langle x_{S}x_{T}\mid S,T\in S(G),S\cap T\neq\emptyset\rangle.$
The ideal $K_{G}$ is called the Kempe ideal of $G$. Then our second main result is the following:
Theorem 1.2.
Let $G$ be a graph on $[d]$ and $<$ a monomial order on $R[G]$, and let $\{{\mathbf{x}}_{f_{1}},\dots,{\mathbf{x}}_{f_{s}}\}$ be the set of all standard monomials of degree $k$ with respect to the initial ideal ${\rm in}_{<}(K_{G})$. Then
$\{f_{1},\dots,f_{s}\}\cap\mathcal{C}_{k}(G)$
is a complete representative system for ${\rm kc}(G,k)$.
As a consequence, the number of $k$-Kempe classes ${\rm Kc}(G,k)$ can be computed by Hilbert functions (Corollary 6.9).
Finally, by using Theorems 1.1 and 1.2 and techniques on Gröbner bases, we introduce several algorithms related to Kempe equivalence. Specifically, our algorithms perform:
  1. (1)
    Determination of Kempe equivalence (Algorithm 7.2);
  2. (2)
    Computation of a complete representative system for ${\rm kc}(G,k)$ (Algorithm 7.3);
  3. (3)
    Enumeration of $k$-colorings that are Kempe equivalent (Algorithm 7.4);
  4. (4)
    Construction of a sequence of Kempe switchings between two Kempe equivalent $k$-colorings (Algorithm 7.6).
The present paper is organized as follows: In Section 2, we will recall the definition of stable set ideals and explain a relationship between stable set ideals and Kempe equivalence. Section 3 will give a brief introduction on Gröbner bases. In Section 4, we will define $2$-coloring ideals and see their algebraic properties. In Section 5, a proof of Theorem 1.1 will be given. In Section 6, we will define Kempe ideals and prove Theorem 1.2. Finally, in Section 7, we will introduce several algorithms related to Kempe equivalence.
2. Stable set ideals
In this section, we define stable set rings and explain a relationship between stable set ideals and Kempe equivalence. Let $G$ be a graph on the vertex set $[d]$ with the edge set $E(G)$. Given a subset $S\subset[d]$, let $G[S]$ denote the induced subgraph of $G$ on the vertex set $S$. A subset $S\subset[d]$ is called a stable set (or an independent set) of $G$ if $\{i,j\}\notin E(G)$ for all $i,j\in S$ with $i\neq j$. Namely, a subset $S\subset[d]$ is stable if and only if $G[S]$ is an empty graph. In particular, the empty set $\emptyset$ and any singleton $\{i\}$ with $i\in[d]$ are stable. Denote $S(G)=\{S_{1},\ldots,S_{n}\}$ the set of all stable sets of $G$. Given a subset $S\subset[d]$, we associate the $(0,1)$-vector $\rho(S)=\sum_{j\in S}{\mathbf{e}}_{j}$. Here ${\mathbf{e}}_{j}$ is the $j$th unit coordinate vector in ${\mathbb{R}}^{d}$. For example, $\rho(\emptyset)=(0,\ldots,0)\in{\mathbb{R}}^{d}$. Let ${\mathbb{K}}[{\mathbf{t}},s]:={\mathbb{K}}[t_{1},\ldots,t_{d},s]$ be the polynomial ring in $d+1$ variables over a field ${\mathbb{K}}$. Given a nonnegative integer vector ${\mathbf{a}}=(a_{1},\ldots,a_{d})\in{\mathbb{Z}}_{\geq 0}^{d}$, we write ${\mathbf{t}}^{{\mathbf{a}}}:=t_{1}^{a_{1}}t_{2}^{a_{2}}\cdots t_{d}^{a_{d}}\in{\mathbb{K}}[{\mathbf{t}},s]$. The stable set ring of $G$ is
${\mathbb{K}}[G]:={\mathbb{K}}[{\mathbf{t}}^{\rho(S_{1})}s,\ldots,{\mathbf{t}}^{\rho(S_{n})}s]\subset{\mathbb{K}}[{\mathbf{t}},s].$
We regard ${\mathbb{K}}[G]$ as a homogeneous algebra by setting each $\deg({\mathbf{t}}^{\rho(S_{i})}s)=1$. Note that ${\mathbb{K}}[G]$ is a toric ring. Let $R[G]={\mathbb{K}}[x_{S_{1}},\ldots,x_{S_{n}}]$ denote the polynomial ring in $n$ variables over ${\mathbb{K}}$ with each $\deg(x_{S_{i}})=1$. The stable set ideal of $G$ is the kernel of the surjective homomorphism $\pi:R[G]\to{\mathbb{K}}[G]$ defined by $\pi(x_{S_{i}})={\mathbf{t}}^{\rho(S_{i})}s$ for $1\leq i\leq n$. Note that $I_{G}$ is a toric ideal, and hence a prime ideal generated by homogeneous binomials. The toric ring ${\mathbb{K}}[G]$ is called quadratic if $I_{G}$ is generated by quadratic binomials. We say that “$I_{G}$ is generated by quadratic binomials” even if $I_{G}=\{0\}$ (or equivalently, $G$ is complete). It is easy to see that a homogeneous binomial $x_{S_{i_{1}}}\cdots x_{S_{i_{r}}}-x_{S_{j_{1}}}\cdots x_{S_{j_{r}}}\in R[G]$ belongs to $I_{G}$ if and only if $\bigcup_{\ell=1}^{r}S_{i_{\ell}}=\bigcup_{\ell=1}^{r}S_{j_{\ell}}$ as multisets. See, e.g., [3] for details on toric rings and toric ideals.
We can describe a system of generators of $I_{G}$ in terms of $k$-colorings. Given a graph $G$ on the vertex set $[d]$, and ${\mathbf{a}}=(a_{1},\ldots,a_{d})\in{\mathbb{Z}}_{\geq 0}^{d}$, let $G_{\mathbf{a}}$ be the graph obtained from $G$ by replacing each vertex $i\in[d]$ with a complete graph $G^{(i)}$ of $a_{i}$ vertices (if $a_{i}=0$, then just delete the vertex $i$), and joining all vertices $x\in G^{(i)}$ and $y\in G^{(j)}$ such that $\{i,j\}$ is an edge of $G$. In particular, if ${\mathbf{a}}=(1,\ldots,1)$, then $G_{\mathbf{a}}=G$. If ${\mathbf{a}}={\bf 0}$, then $G_{\mathbf{a}}$ is the null graph (a graph without vertices). In addition, if ${\mathbf{a}}$ is a $(0,1)$-vector, namely, ${\mathbf{a}}\in\{0,1\}^{d}$, then $G_{\mathbf{a}}$ is an induced subgraph of $G$. If ${\mathbf{a}}$ is a positive vector, then $G_{\mathbf{a}}$ is called a replication graph of $G$. In general, $G_{\mathbf{a}}$ is a replication graph of an induced subgraph of $G$. Given a $k$-coloring $f$ of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$, we associate $f$ with a monomial
${\mathbf{x}}_{f}:=x_{S_{i_{1}}}\cdots x_{S_{i_{k}}}\in R[G],$
where $S_{i_{\ell}}=\{j\in[d]\mid G^{(j)}\cap f^{-1}(\ell)\neq\emptyset\}$ for $\ell=1,2,\ldots,k$. Conversely, let $m=x_{S_{i_{1}}}\cdots x_{S_{i_{k}}}\in R[G]$ be a monomial of degree $k$. Then, for ${\mathbf{a}}=(a_{1},\ldots,a_{n})$ with $a_{p}=|\{\ell:p\in S_{i_{\ell}}\}|$, there exists a $k$-coloring $f$ of $G_{{\mathbf{a}}}$ such that ${\mathbf{x}}_{f}=m$ (see [7, Lemma 3.2]). Note that, for $k$-colorings $f$ and $g$ of an induced subgraph of $G$, ${\mathbf{x}}_{f}={\mathbf{x}}_{g}$ if and only if $g$ is obtained from $f$ by permuting colors. It is easy to see that $f\sim_{k}g$ if $g$ is obtained from $f$ by permuting colors. In this paper, we identify $f$ and $g$ if $g$ is obtained from $f$ by permuting colors. Then we can describe a system of generators of $I_{G}$ as follows:
Proposition 2.1 ([7, Theorem 3.3]).
Let $G$ be a graph on $[d]$. Then one has
${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in I_{G}\iff\mbox{$f$ and $g$ are $k$-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and $k\geq\chi(G_{{\mathbf{a}}})$}$
and
$I_{G}$ $=\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are $k$-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and $k\geq\chi(G_{{\mathbf{a}}})$}\rangle$
$=\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are k-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in\{0,1,\dots,k\}^{d}$ and $k\geq\chi(G_{{\mathbf{a}}})$}\rangle.$
Now, we explain a relationship between $I_{G}$ and Kempe equivalence.
Proposition 2.2 ([7, Theorem 1.3]).
Let $G$ be a graph on $[d]$. Then $I_{G}$ is generated by quadratic binomials if and only if for any ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and $k\geq\chi(G_{{\mathbf{a}}})$, all $k$-colorings of $G_{{\mathbf{a}}}$ are Kempe equivalent, namely, one has
${\rm Kc}(G_{{\mathbf{a}}},k)=1.$
We consider the ideal $\langle[I_{G}]_{2}\rangle$ of $R[G]$ generated by all quadratic binomials of $I_{G}$, namely,
$\langle[I_{G}]_{2}\rangle$ $=\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are 2-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$}\rangle$
$=\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are 2-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in\{0,1,2\}^{d}$}\rangle\subset R[G].$
Proposition 2.3.
Let $G$ be a graph on $[d]$ and $f,g$ $k$-colorings of $G$. Then $f\sim_{k}g$ if and only if ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in\langle[I_{G}]_{2}\rangle$.
We will see a proof of this proposition in Section 5 (Theorem 5.1) by using a similar discussion in the proof of Proposition 2.2 in [7].
3. Gröbner bases
In this section, we give a brief introduction on Gröbner bases. Let $R={\mathbb{K}}[x_{1},x_{2},\ldots,x_{n}]$ be a polynomial ring over a field ${\mathbb{K}}$ with $\deg(x_{i})=1$, and denote $\mathcal{M}_{n}$ the set of all monomials in the variables $x_{1},\ldots,x_{n}$. A monomial order on $R$ is a total order $<$ on $\mathcal{M}_{n}$ such that
  1. (1)
    $1<u$ for all $1\neq u\in\mathcal{M}_{n}$;
  2. (2)
    if $u,v\in\mathcal{M}_{n}$ and $u<v$, then $uw<vw$ for all $w\in\mathcal{M}_{n}$.
We give an example of a monomial order.
Example 3.1.
Let $u=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$ and $v=x_{1}^{b_{1}}\cdots x_{n}^{b_{n}}$ be monomials in $\mathcal{M}_{n}$. We define the total order $<_{\rm rev}$ on $\mathcal{M}_{n}$ by setting $u<_{\rm rev}v$ if either (i) $\sum_{i=1}^{n}a_{i}<\sum_{i=1}^{n}b_{i}$, or (ii) $\sum_{i=1}^{n}a_{i}=\sum_{i=1}^{n}b_{i}$ and the rightmost nonzero component of the vector $(b_{1}-a_{1},b_{2}-a_{2},\ldots,b_{n}-a_{n})$ is negative. It then follows that $<_{\rm rev}$ is a monomial order on $R$, which is called the (graded) reverse lexicographic order on $R$ induced by the ordering $x_{1}>x_{2}>\cdots>x_{n}$. By reordering the variables, we can obtain another reverse lexicographic order on $R$. Hence, there are $n!$ reverse lexicographic orders on $R$.
Fix a monomial order $<$ on $R$. For a nonzero polynomial $f$ of $R$, the support of $f$, denoted by ${\rm supp}(f)$, is the set of all monomials appearing in $f$ and the initial monomial ${\rm in}_{<}(f)$ of $f$ with respect to $<$ is the largest monomial belonging to ${\rm supp}(f)$ with respect to $<$. Let $I$ be a nonzero ideal of $R$. Then the initial ideal ${\rm in}_{<}(I)$ of $I$ with respect to $<$ is defined as follows:
${\rm in}_{<}(I)=\langle{\rm in}_{<}(f)\mid 0\neq f\in I\rangle\subset R.$
In general, even if $I=\langle f_{1},\ldots,f_{s}\rangle$, it is not necessarily true that ${\rm in}_{<}(I)=\langle{\rm in}_{<}(f_{1}),\ldots,{\rm in}_{<}(f_{s})\rangle$. A finite set $\{g_{1},\ldots,g_{s}\}$ of nonzero polynomials belonging to $I$ is called a Gröbner basis of $I$ with respect to $<$ if ${\rm in}_{<}(I)=\langle{\rm in}_{<}(g_{1}),\ldots,{\rm in}_{<}(g_{s})\rangle$. Note that if $\{g_{1},\ldots,g_{s}\}$ is a Gröbner basis of $I$, then $I=\langle g_{1},\ldots,g_{s}\rangle$. It is known that $I$ always has a Gröbner basis. A Gröbner basis ${\mathcal{G}}$ for $I$ with respect to $<$ is called reduced if the following conditions hold:
  1. (1)
    for all $p\in{\mathcal{G}}$, the coefficient of ${\rm in}_{<}(p)$ in $p$ equals $1$;
  2. (2)
    for all $p\in{\mathcal{G}}$, no monomial of $p$ lies in $\langle{\rm in}_{<}(g)\mid g\in{\mathcal{G}}\setminus\{p\}\rangle$.
A nonzero ideal $I$ has a unique reduced Gröbner basis with respect to a fixed monomial order $<$. In particular, given a Gröbner basis, we can easily get the reduced Gröbner basis from it.
Next we introduce a method to determine if a finite system of generators of $I$ is a Gröbner basis of $I$ with respect to $<$. For two nonzero polynomials $f$ and $g$ in $R$, the polynomial
$S(f,g)=\frac{m}{c_{f}\cdot{\rm in}_{<}(f)}\cdot f-\frac{m}{c_{g}\cdot{\rm in}_{<}(g)}\cdot g$
is called the $S$-polynomial of $f$ and $g$, where $c_{f}$ is the coefficient of ${\rm in}_{<}(f)$ in $f$ and $c_{g}$ is that of ${\rm in}_{<}(g)$ in $g$, and $m$ is the least common multiple of the initial monomials ${\rm in}_{<}(f)$ and ${\rm in}_{<}(g)$.
Lemma 3.2 (Buchberger Criterion [2, Chapter 2, §9, Theorem 3]).
Let $I$ be a nonzero ideal of $R$, $\mathcal{G}=\{g_{1},\ldots,g_{s}\}$ a finite system of generators of $I$. Then ${\mathcal{G}}$ is a Gröbner basis of $I$ with respect to a monomial order $<$ on $R$ if and only if the remainder of $S$-polynomial $S(g_{i},g_{j})$ on division by ${\mathcal{G}}$ is 00 for all $i\neq j$.
The algorithm called Buchberger Algorithm ([2, Chapter 2, §7, Theorem 2]) is based on Buchberger Criterion, and computes a Gröbner basis for $I$ from a finite system of generators of $I$.
As applications of Gröbner bases, we can determine if a polynomial $f$ belongs to an ideal. This problem is called the ideal membership problem. If ${\mathcal{G}}$ is a Gröbner basis for an ideal $I$ of $R$ with respect to a monomial order $<$, then every polynomial of $f\in R$ has a unique remainder on division by ${\mathcal{G}}$. The remainder is called the normal form of $f$ with respect to ${\mathcal{G}}$.
Lemma 3.3 ([2, Chapter 2, §6, Corollary 2]).
Let $I$ be a nonzero ideal of $R$ and $\mathcal{G}$ a Gröbner basis of $I$ with respect to a monomial order $<$ on $R$. Fix a polynomial $f$ in $R$. Then $f\in I$ if and only if the normal form of $f$ with respect to ${\mathcal{G}}$ equals 00.
Next, we review a method to compute Hilbert functions. Let $I$ be a graded ideal of $R$. The numerical function $H(R/I,-):{\mathbb{Z}}_{\geq 0}\to{\mathbb{Z}}_{\geq 0}$ with $H(R/I,k)=\dim_{{\mathbb{K}}}R_{k}/I_{k}$ is called the Hilbert function of $R/I$, where $R_{k}$ (resp. $I_{k}$) is the homogeneous part of degree $k$ of $R$ (resp. $I$) and $\dim_{{\mathbb{K}}}R_{k}/I_{k}$ is the dimension of $R_{k}/I_{k}$ as ${\mathbb{K}}$-vector spaces. For a monomial order $<$ on $R$, a monomial $u\in\mathcal{M}_{n}$ is said to be standard with respect to ${\rm in}_{<}(I)$ if $u\notin{\rm in}_{<}(I)$.
Lemma 3.4 ([3, Theorem 1.19]).
Let $I$ be a nonzero graded ideal of $R$ and fix a monomial order $<$ on $R$. Let $\mathcal{B}_{k}$ denote the set of standard monomials of degree $k$ with respect to ${\rm in}_{<}(I)$. Then $\mathcal{B}_{k}$ is a ${\mathbb{K}}$-basis of $S_{k}/I_{k}$ as a ${\mathbb{K}}$-vector space. In particular, one has
$H(R/I,k)=|\mathcal{B}_{k}|.$
Finally, we recall a property of the saturation of an ideal. For a nonzero ideal $I$ of $R$ and a polynomial $f\in R$, the saturation of $I$ with respect to $f$ is the ideal
$I:f^{\infty}=\{g\in R\mid\mbox{there exists $i>0$ such that $f^{i}g\in I$}\}.$
Lemma 3.5 ([3, Proposition 1.40]).
Let $I$ be a nonzero graded ideal of $R$ and let $\mathcal{G}$ be the reduced Gröbner basis of $I$ with respect to the reverse lexicographic order induced by $x_{1}>x_{2}>\cdots>x_{n}$. Then
[…]
is a Gröbner basis of  $I:x_{n}^{\infty}$.
4. 2-coloring ideals
Given a graph $G$ on $[d]$, we define the following two ideals of $R[G]$:
$J_{G}:=$ $\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are$2$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in\{0,1\}^{d}$}\rangle$
$=$ $\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are$2$-colorings of an induced subgraph of $G$}\rangle,$
$L_{G}:=$ $\langle x_{S\setminus\{i\}}x_{\{i\}}-x_{S}x_{\emptyset}\mid i\in S\in S(G),\ |S|\geq 2\rangle.$
We call $J_{G}$ the $2$-coloring ideal of $G$. Note that inclusions of ideals
(4.1) $L_{G}\subset J_{G}\subset\langle[I_{G}]_{2}\rangle\subset I_{G}$
hold. Moreover, from Proposition 2.1, for $I\in\{\langle[I_{G}]_{2}\rangle,J_{G},L_{G}\}$, if ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in I$, then $f$ and $g$ are $k$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$.
In this section, we discuss some properties of these ideals.
Lemma 4.1.
Let $G$ be a graph on $[d]$. Then one has $L_{G}:{\mathbf{x}}_{\emptyset}^{\infty}=J_{G}:{\mathbf{x}}_{\emptyset}^{\infty}=\langle[I_{G}]_{2}\rangle:{\mathbf{x}}_{\emptyset}^{\infty}=I_{G}$.
Proof.
Since $L_{G}\subset J_{G}\subset\langle[I_{G}]_{2}\rangle\subset I_{G}$, we have $L_{G}:{\mathbf{x}}_{\emptyset}^{\infty}\subset J_{G}:{\mathbf{x}}_{\emptyset}^{\infty}\subset\langle[I_{G}]_{2}\rangle:{\mathbf{x}}_{\emptyset}^{\infty}\subset I_{G}:{\mathbf{x}}_{\emptyset}^{\infty}.$ In addition, since $I_{G}$ is prime and does not contain $x_{\emptyset}^{k}$ for any $k$, we have $I_{G}:{\mathbf{x}}_{\emptyset}^{\infty}=I_{G}$. Thus it is sufficient to show that $I_{G}\subset L_{G}:{\mathbf{x}}_{\emptyset}^{\infty}$. Let
$F=\prod_{i=1}^{s}x_{S_{i}}-\prod_{i=1}^{s}x_{S_{i}^{\prime}}\in I_{G},$
where $S_{i}=\{k_{1}^{(i)},\dots,k_{p_{i}}^{(i)}\}$, $S_{i}^{\prime}=\{\ell_{1}^{(i)},\dots,\ell_{q_{i}}^{(i)}\}\in S(G)$ for each $i$. Then
$x_{S_{i}}x_{\emptyset}^{p_{i}-1}-\prod_{j=1}^{p_{i}}x_{\{k_{j}^{(i)}\}}$
$=$ $x_{\emptyset}^{p_{i}-2}\left(x_{S_{i}}x_{\emptyset}-x_{S_{i}\setminus\{k_{1}^{(i)}\}}x_{\{k_{1}^{(i)}\}}\right)+x_{\emptyset}^{p_{i}-3}x_{\{k_{1}^{(i)}\}}\left(x_{S_{i}\setminus\{k_{1}^{(i)}\}}x_{\emptyset}-x_{S_{i}\setminus\{k_{1}^{(i)},k_{2}^{(i)}\}}x_{\{k_{2}^{(i)}\}}\right)$
$+\cdots+\left(\prod_{j=1}^{p_{i}-2}x_{\{k_{j}^{(i)}\}}\right)\left(x_{\{k_{p_{i}-1}^{(i)},k_{p_{i}}^{(i)}\}}x_{\emptyset}-x_{\{k_{p_{i}}^{(i)}\}}x_{\{k_{p_{i}-1}^{(i)}\}}\right)$
belongs to $L_{G}$. Note that if binomials $u_{1}-v_{1},\dots,u_{s}-v_{s}$ belong to $L_{G}$, then
$u_{1}\dots u_{s}-v_{1}\dots v_{s}$
$=$ $u_{2}\cdots u_{s}(u_{1}-v_{1})+u_{3}\cdots u_{s}v_{1}(u_{2}-v_{2})+u_{4}\cdots u_{s}v_{1}v_{2}(u_{3}-v_{3})+\cdots+v_{1}\cdots v_{s-1}(u_{s}-v_{s})$
belongs to $L_{G}$. Hence
$G_{1}=x_{\emptyset}^{\sum_{i=1}^{s}(p_{i}-1)}\prod_{i=1}^{s}x_{S_{i}}-\prod_{i=1}^{s}\prod_{j=1}^{p_{i}}x_{\{k_{j}^{(i)}\}}$
belongs to $L_{G}$. By the same argument,
$G_{2}=x_{\emptyset}^{\sum_{i=1}^{s}(q_{i}-1)}\prod_{i=1}^{s}x_{S_{i}^{\prime}}-\prod_{i=1}^{s}\prod_{j=1}^{q_{i}}x_{\{\ell_{j}^{(i)}\}}$
belongs to $L_{G}$. Since $F$ belongs to $I_{G}$, we have
$\sum_{i=1}^{s}(p_{i}-1)=-s+\sum_{i=1}^{s}p_{i}=-s+\sum_{i=1}^{s}q_{i}=\sum_{i=1}^{s}(q_{i}-1),$
and
$\prod_{i=1}^{s}\prod_{j=1}^{p_{i}}x_{\{k_{j}^{(i)}\}}=\prod_{i=1}^{s}\prod_{j=1}^{q_{i}}x_{\{\ell_{j}^{(i)}\}}.$
Thus
$x_{\emptyset}^{\sum_{i=1}^{s}(p_{i}-1)}F=G_{1}-G_{2}$
belongs to $L_{G}$. This implies that $F\in L_{G}:x^{\infty}_{\emptyset}$. Hence one has $I_{G}\subset L_{G}:x^{\infty}_{\emptyset}$. ∎
Therefore, combining Lemmata 3.5 and 4.1, we obtain the following proposition.
Proposition 4.2.
Let $G$ be a graph on $[d]$ and let $I\in\{\langle[I_{G}]_{2}\rangle,J_{G},L_{G}\}$. If $\mathcal{G}$ is the reduced Gröbner basis of $I$ with respect to a reverse lexicographic order such that $x_{S}\geq x_{\emptyset}$ for any $S\in S(G)$, then
[…]
is a Gröbner basis of $I_{G}$.
Next, we discuss when equality holds in the inclusions $L_{G}\subset J_{G}\subset\langle[I_{G}]_{2}\rangle\subset I_{G}$. In particular, we characterize when these ideals are prime. Let $\overline{G}$ denote the complement graph of a graph $G$.
Proposition 4.3.
Let $G$ be a graph on $[d]$. Then one has the following.
  1. (a)
    Let $I\in\{\langle[I_{G}]_{2}\rangle,J_{G},L_{G}\}$. Then $I$ is prime $\iff$ $I=I_{G}$.
  2. (b)
    $I_{G}=\langle[I_{G}]_{2}\rangle\iff I_{G}$ is generated by quadratic binomials.
  3. (c)
    $\langle[I_{G}]_{2}\rangle=J_{G}\iff\overline{G}$ has no $3$-cycles.
  4. (d)
    $J_{G}=L_{G}\iff G$ is a complete multipartite graph on the vertex set $V_{1}\sqcup\cdots\sqcup V_{t}$ with $|V_{j}|\leq 3$ for each $j$.
Proof.
(a) Let $I\in\{\langle[I_{G}]_{2}\rangle,J_{G},L_{G}\}$. Recall that the toric ideal $I_{G}$ is prime. Hence $I$ is prime if $I_{G}=I$. Conversely, suppose that $I$ is prime. Since $x^{k}_{\emptyset}\notin I$ holds for any $k$, it follows that $I=I:x_{\emptyset}^{\infty}$. From Lemma 4.1, we have $I=I:x_{\emptyset}^{\infty}=I_{G}$.
(b) Trivial.
(c) If $\overline{G}$ has a $3$-cycle $(i_{1},i_{2},i_{3})$, then $x_{\{i_{1},i_{2},i_{3}\}}x_{\{i_{1}\}}-x_{\{i_{1},i_{2}\}}x_{\{i_{1},i_{3}\}}\in\langle[I_{G}]_{2}\rangle\setminus J_{G}$. Suppose that $\overline{G}$ has no $3$-cycles. Then $|S|\leq 2$ for all $S\in S(G)$. Let $h=x_{S_{1}}x_{S_{2}}-x_{S_{3}}x_{S_{4}}\in\langle[I_{G}]_{2}\rangle$. Since $h$ belongs to $I_{G}$, $S_{1}\cup S_{2}$ coincides with $S_{3}\cup S_{4}$ as multisets, and in particular, we have $S_{1}\cap S_{2}=S_{3}\cap S_{4}$. Suppose that $S_{1}\cap S_{2}$ is not empty. Let $i\in S_{1}\cap S_{2}$. Then $(S_{1}\setminus\{i\})\cup(S_{2}\setminus\{i\})$ coincides with $(S_{3}\setminus\{i\})\cup(S_{4}\setminus\{i\})$ as multisets. Since $|S_{j}\setminus\{i\}|\leq 1$ for each $j=1,2,3,4$, it follows that $(S_{1},S_{2})$ is equal to either $(S_{3},S_{4})$ or $(S_{4},S_{3})$. Then $h=0$, a contradiction. Thus $S_{1}\cap S_{2}=\emptyset$, and hence $h\in J_{G}$. Therefore we have $\langle[I_{G}]_{2}\rangle\subset J_{G}$, and hence $\langle[I_{G}]_{2}\rangle=J_{G}$.
(d) Suppose that $J_{G}=L_{G}$. If $\overline{G}$ has a path of length two $P_{3}=(i,j,k)$ as an induced subgraph, then it follows that $x_{\{i,j\}}x_{\{k\}}-x_{\{i\}}x_{\{j,k\}}\in J_{G}\setminus L_{G}$, a contradiction. Hence $\overline{G}$ has no $P_{3}$ as an induced subgraph. It is known that $\overline{G}$ has no $P_{3}$ as an induced subgraph if and only if $G$ is a complete multipartite graph. Suppose that $G$ has a part $V_{\alpha}$ with $|V_{\alpha}|\geq 4$. It then follows that $x_{\{i,j\}}x_{\{k,l\}}-x_{\{i,k\}}x_{\{j,l\}}\in J_{G}\setminus L_{G}$ where $i,j,k,l$ are distinct vertices in $V_{\alpha}$, a contradiction. Hence $G$ is a complete multipartite graph on the vertex set $V_{1}\sqcup\cdots\sqcup V_{t}$ with $|V_{j}|\leq 3$ for each $j$.
Suppose that $G$ is a complete multipartite graph on the vertex set $V_{1}\sqcup\cdots\sqcup V_{t}$ with $|V_{j}|\leq 3$ for each $j$. Then each $S\in S(G)$ is a subset of $V_{j}$ for some $j$. It is enough to show that $J_{G}\subset L_{G}$. Let $h=x_{S_{1}}x_{S_{2}}-x_{S_{3}}x_{S_{4}}$ be a nonzero binomial in $J_{G}$. Then $S_{1}\cap S_{2}=S_{3}\cap S_{4}=\emptyset$ and $S_{1}\cup S_{2}=S_{3}\cup S_{4}$. Suppose that $S_{i}\subset V_{j_{i}}$ for each $i=1,2,3,4$.
Case 1. ($S_{1}=\emptyset$.) Then $h=x_{\emptyset}x_{S_{2}}-x_{S_{3}}x_{S_{4}}$, where $|S_{3}|,|S_{4}|\geq 1$ and $|S_{3}|+|S_{4}|=|S_{2}|\leq 3$. Since either $|S_{3}|$ or $|S_{4}|$ equals one, $h\in L_{G}$.
Case 2. ($S_{j}\neq\emptyset$ for each $i$.) Since $S_{1}\cup S_{2}=S_{3}\cup S_{4}$, we may assume that $j_{1}=j_{3}$ and $j_{2}=j_{4}$. If $j_{1}\neq j_{2}$, then $S_{1}=S_{3}$ and $S_{2}=S_{4}$, and hence $h=0$. This is a contradiction. Thus we have $j_{1}=j_{2}=j_{3}=j_{4}$. Since $2\leq|S_{1}|+|S_{2}|=|S_{3}|+|S_{4}|\leq 3$, we may assume that $|S_{1}|=|S_{3}|=1$. Then $h=(x_{S_{1}}x_{S_{2}}-x_{\emptyset}x_{S_{1}\cup S_{2}})+(x_{\emptyset}x_{S_{3}\cup S_{4}}-x_{S_{3}}x_{S_{4}})$ belongs to $L_{G}$. ∎
Now, we return to the inequalities (4.1). If $G$ is a complete multipartite graph, then $I_{G}$ is generated by quadratic binomials ([6, Theorem 3.1]). Hence from Proposition 4.3 we know that $L_{G}=J_{G}$ implies $\langle[I_{G}]_{2}\rangle=I_{G}$. Therefore, there are the following $6$ cases:
  1. (i)
    $L_{G}=J_{G}=\langle[I_{G}]_{2}\rangle=I_{G}$.
  2. (ii)
    $L_{G}=J_{G}\subsetneq\langle[I_{G}]_{2}\rangle=I_{G}$.
  3. (iii)
    $L_{G}\subsetneq J_{G}=\langle[I_{G}]_{2}\rangle=I_{G}$.
  4. (iv)
    $L_{G}\subsetneq J_{G}=\langle[I_{G}]_{2}\rangle\subsetneq I_{G}$.
  5. (v)
    $L_{G}\subsetneq J_{G}\subsetneq\langle[I_{G}]_{2}\rangle=I_{G}$.
  6. (vi)
    $L_{G}\subsetneq J_{G}\subsetneq\langle[I_{G}]_{2}\rangle\subsetneq I_{G}$.
We give an example for each case.
Example 4.4.
First, we recall that for any bipartite graph $G$, $I_{G}$ is generated by quadratic binomials.
A complete graph $K_{d}$ of $d$ vertices satisfies the condition (i). In fact, one has $I_{K_{d}}=\{0\}$. For a $(3,3)$-complete bipartite graph $K_{3,3}$, since $\overline{K_{3,3}}$ has a $3$-cycle, the condition (ii) holds. For a path of length three $P_{4}$, since $\overline{P_{4}}$ is also a path of length three, the condition (iii) holds. For $G=\overline{C_{6}}$, $I_{G}$ is not generated by quadratic binomials from [4, Proposition 11]. Since $\overline{G}=C_{6}$ has no $3$-cycles and $G$ is not a complete multipartite graph, the condition (iv) is satisfied. For a $6$-cycle $C_{6}$, since $\overline{C_{6}}$ has a $3$-cycle and $C_{6}$ is not a complete multipartite graph, the condition (v) holds. Finally, we consider the following graph $G$:
It then follows from [6, Theorem 1.7] that $I_{G}$ is not generated by quadratic binomials. Since $\overline{G}$ has a $3$-cycle and $G$ is not a complete multipartite graph, the condition (vi) holds.
5. Examining Kempe equivalence
In this section, we prove Theorem 1.1 and Proposition 2.3. In fact, we show the following.
Theorem 5.1.
Let $G$ be a graph on $[d]$ and let $f$ and $g$ be $k$-colorings of an induced subgraph of $G$. Then the following conditions are equivalent:
  1. (i)
    $f\sim_{k}g$;
  2. (ii)
    ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in\langle[I_{G}]_{2}\rangle$;
  3. (iii)
    ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$.
Proof.
Suppose that $f$ and $g$ are $k$-colorings of an induced subgraph $G_{0}$ of $G$. Since $J_{G}\subset\langle[I_{G}]_{2}\rangle$ holds, we have (iii) $\Rightarrow$ (ii).
(i) $\Rightarrow$ (iii). Suppose that $f\sim_{k}g$ and ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\notin J_{G}$. Let $f_{0},f_{1},\ldots,f_{s}$ be a sequence of $k$-colorings of $G_{0}$ such that $f_{0}=f$, $f_{s}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching. We may assume that $s\geq 1$ is minimal among all $k$-colorings $f$ and $g$ such that $f\sim_{k}g$ and ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\notin J_{G}$. Suppose that the Kempe switching from $f$ to $f_{1}$ is obtained by a connected component $H$ of the induced subgraph $G_{0}[f^{-1}(\mu)\cup f^{-1}(\eta)]$ by setting
$f_{1}(x)=\begin{cases}f(x)&x\notin H,\\ \mu&x\in H\mbox{ and }f(x)=\eta,\\ \eta&x\in H\mbox{ and }f(x)=\mu.\\ \end{cases}$
Let $f^{\prime}=f|_{G^{\prime}}$ and $f_{1}^{\prime}=f_{1}|_{G^{\prime}}$ be the restrictions of $f$ and $f_{1}$ to $G^{\prime}:=G_{0}[f^{-1}(\mu)\cup f^{-1}(\eta)]$, respectively. Since $f^{\prime}$ and $f_{1}^{\prime}$ are 2-colorings of ${G^{\prime}}$, ${\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{f_{1}^{\prime}}$ belongs to $J_{G}$. Then
${\mathbf{x}}_{f}-{\mathbf{x}}_{g}=\frac{{\mathbf{x}}_{f}}{{\mathbf{x}}_{f^{\prime}}}({\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{f_{1}^{\prime}})+\left(\frac{{\mathbf{x}}_{f}}{{\mathbf{x}}_{f^{\prime}}}{\mathbf{x}}_{f_{1}^{\prime}}-{\mathbf{x}}_{g}\right)=\frac{{\mathbf{x}}_{f}}{{\mathbf{x}}_{f^{\prime}}}({\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{f_{1}^{\prime}})+{\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g}.$
If $s=1$, then $f_{1}=g$ and hence
${\mathbf{x}}_{f}-{\mathbf{x}}_{g}=\frac{{\mathbf{x}}_{f}}{{\mathbf{x}}_{f^{\prime}}}({\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{f_{1}^{\prime}})\in J_{G}.$
We may assume that $s\geq 2$. By the hypothesis on $s$, ${\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g}$ belongs to $J_{G}$. Hence ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ belongs to $J_{G}$, a contradiction.
(ii) $\Rightarrow$ (i). Suppose that $f\not\sim_{k}g$ and ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in\langle[I_{G}]_{2}\rangle$. Since $\langle[I_{G}]_{2}\rangle$ is a binomial ideal, it then follows from [3, Lemma 3.8] that there exists an expression
(5.1) ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}=\sum_{r=1}^{s}{\bf x}_{w_{r}}({\mathbf{x}}_{f_{r}}-{\mathbf{x}}_{g_{r}}),$
where $f_{r}$ and $g_{r}$ are 2-colorings of $G_{{\mathbf{a}}_{r}}$ with ${\mathbf{a}}_{r}\in{\mathbb{Z}}_{\geq 0}^{d}$ for each $r$. We may assume that $s\geq 1$ is minimal among such $f$ and $g$ with $f\not\sim_{k}g$ and ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in\langle[I_{G}]_{2}\rangle$. Since ${\mathbf{x}}_{f}$ must appear in the right-hand side of (5.1), we may assume that ${\mathbf{x}}_{f}={\mathbf{x}}_{w_{1}}{\mathbf{x}}_{f_{1}}$. Then ${\mathbf{x}}_{f}$ is divided by ${\mathbf{x}}_{f_{1}}$, and $f_{1}$ is the restriction of $f$ to $G^{\prime}=G_{0}[f^{-1}(\mu)\cup f^{-1}(\nu)]$ for some $\mu$ and $\nu$. Since $G^{\prime}$ has a 2-coloring, it is a bipartite graph. Then $f_{1}$ and $g_{1}$ are Kempe equivalent. Let $f^{\prime}$ be a coloring of $G_{0}$ defined by
$f^{\prime}(x)=\begin{cases}g_{1}(x)&f(x)\in\{\mu,\nu\},\\ f(x)&\mbox{otherwise.}\end{cases}$
Then one obtains $f\sim_{k}f^{\prime}$. Moreover, we have
(5.2) ${\mathbf{x}}_{f}-{\bf x}_{w_{1}}({\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g_{1}})={\mathbf{x}}_{f^{\prime}}.$
If $s=1$, then
${\mathbf{x}}_{g}={\mathbf{x}}_{f}-{\bf x}_{w_{1}}({\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g_{1}})={\mathbf{x}}_{f^{\prime}},$
and hence $f\sim_{k}f^{\prime}=g$, a contradiction. We may assume that $s\geq 2$. Since
${\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{g}={\mathbf{x}}_{f}-{\bf x}_{w_{1}}({\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g_{1}})-{\mathbf{x}}_{g}=\sum_{r=2}^{s}{\bf x}_{w_{r}}({\mathbf{x}}_{f_{r}}-{\mathbf{x}}_{g_{r}})\in\langle[I_{G}]_{2}\rangle,$
$f^{\prime}\sim_{k}g$ by the hypothesis on $s$. Thus $f\sim_{k}f^{\prime}\sim_{k}g$, a contradiction. ∎
We see examples of Theorem 5.1.
Example 5.2.
Let $G$ be the graph as follows:
111122223333444455556666
We consider two $3$-colorings $f$ and $g$ of $G$ defined by
$f(i)=\begin{cases}1&i\in\{1,5\},\\ 2&i\in\{2,6\},\\ 3&i\in\{3,4\}\end{cases}\mbox{ and }g(i)=\begin{cases}1&i\in\{1,3,5\},\\ 2&i\in\{2,6\},\\ 3&i=4.\end{cases}$
Then one has
${\mathbf{x}}_{f}-{\mathbf{x}}_{g}=x_{\{1,5\}}x_{\{2,6\}}x_{\{3,4\}}-x_{\{1,3,5\}}x_{\{2,6\}}x_{\{4\}}=x_{\{2,6\}}(x_{\{1,5\}}x_{\{3,4\}}-x_{\{1,3,5\}}x_{\{4\}}).$
Since $x_{\{1,5\}}x_{\{3,4\}}-x_{\{1,3,5\}}x_{\{4\}}\in J_{G}$, it then follows from Theorem 5.1 that $f\sim_{3}g$.
Example 5.3.
Let $G$ be the graph as follows:
111122223333444455556666
We consider two $3$-colorings $f$ and $g$ of $G$ defined by
$f(i)=\begin{cases}1&i\in\{1,5\},\\ 2&i\in\{2,6\},\\ 3&i\in\{3,4\}\end{cases}\mbox{ and }g(i)=\begin{cases}1&i\in\{1,6\},\\ 2&i\in\{2,4\},\\ 3&i\in\{3,5\}.\end{cases}$
Then one has
${\mathbf{x}}_{f}-{\mathbf{x}}_{g}=x_{\{1,5\}}x_{\{2,6\}}x_{\{3,4\}}-x_{\{1,6\}}x_{\{2,4\}}x_{\{3,5\}}.$
Since ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\notin J_{G}$, it then follows from Theorem 5.1 that $f\not\sim_{3}g$.
See Section 7 for an algorithm to determine if $f\sim_{k}g$ by using the techniques on Gröbner bases.
6. Computing Kempe classes via Kempe ideals
In this section, we give a way to find all $k$-colorings of a graph $G$ up to Kempe equivalence via commutative algebra. In particular, we prove Theorem 1.2. For a graph $G$ on $[d]$, we define the ideal $K_{G}$ of $R$ by
$K_{G}:=J_{G}+M_{G}\subset R[G],$
where
$M_{G}:=\langle x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\rangle\subset R[G].$
The ideal $K_{G}$ is called the Kempe ideal of $G$.
The set of all monomials in $M_{G}$ consists of all monomials associated to $k$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$ such that $G_{{\mathbf{a}}}$ is not an induced subgraph of $G$.
Lemma 6.1.
Let $G$ be a graph on $[d]$. Then ${\mathbf{x}}_{f}=x_{S_{1}}x_{S_{2}}\cdots x_{S_{k}}\notin M_{G}$ if and only if $f$ is a $k$-coloring of an induced subgraph of $G$.
Proof.
Suppose that $f$ is not a $k$-coloring of any induced subgraph of $G$. Then $f$ is a $k$-coloring of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$ such that $a_{j}\geq 2$ for some $j$. Hence there exist integers $1\leq a<b\leq k$ with $i\in S_{a}$ and $i\in S_{b}$. This implies that ${\mathbf{x}}_{f}$ can be divided by a monomial $x_{S_{a}}x_{S_{b}}\in M_{G}$, and hence ${\mathbf{x}}_{f}\in M_{G}$.
Suppose that ${\mathbf{x}}_{f}\in M_{G}$. There exist $S,T\in S(G)$ such that $S\cap T\neq\emptyset$ and $x_{S}x_{T}$ divides ${\mathbf{x}}_{f}$. Let $i\in S\cap T$. Then $i$ appears at least twice in $S_{1}\cup\dots\cup S_{k}$. Thus $f$ is not a $k$-coloring of any induced subgraph of $G$. ∎
As an easy consequence of this lemma, we have the following.
Corollary 6.2.
Let $G$ be a graph on $[d]$. Then one has
$K_{G}=\langle[I_{G}]_{2}\rangle+M_{G}.$
The following lemma is useful to check if a homogeneous binomial in $I_{G}$ belongs to $M_{G}$.
Lemma 6.3.
Let $G$ be a graph on $[d]$ and let ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in I_{G}$. Then the following conditions are equivalent:
  • (i)
    ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in M_{G}$;
  • (ii)
    both ${\mathbf{x}}_{f}$ and ${\mathbf{x}}_{g}$ belong to $M_{G}$;
  • (iii)
    at least one of ${\mathbf{x}}_{f}$ and ${\mathbf{x}}_{g}$ belongs to $M_{G}$.
Proof.
First, (ii) $\Longrightarrow$ (iii) is trivial. Since $M_{G}$ is a monomial ideal, the equivalence (i) $\iff$ (ii) holds. We show that (iii) $\Longrightarrow$ (ii). Suppose that ${\mathbf{x}}_{f}\in M_{G}$, i.e., ${\mathbf{x}}_{f}$ is divided by $x_{S}x_{T}$ where $S,T\in S(G)$ and $S\cap T\neq\emptyset$. Let $i\in S\cap T$. Then $\pi({\mathbf{x}}_{f})$ is divided by $t_{i}^{2}$ since ${\mathbf{x}}_{f}$ is divided by $x_{S}x_{T}$. From ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in I_{G}$, we have $\pi({\mathbf{x}}_{f})=\pi({\mathbf{x}}_{g})$. Hence if ${\mathbf{x}}_{g}=x_{S_{1}}x_{S_{2}}\cdots x_{S_{k}}$, then there exist $1\leq a<b\leq k$ such that $i\in S_{a}\cap S_{b}$. Namely, ${\mathbf{x}}_{g}$ is divided by a monomial $x_{S_{a}}x_{S_{b}}$ such that $i\in S_{a}\cap S_{b}$. Thus ${\mathbf{x}}_{g}\in M_{G}$. ∎
Next, we discuss a Gröbner basis of $K_{G}$.
Proposition 6.4.
Let $G$ be a graph on $[d]$, and let $\mathcal{G}_{1}$ be the reduced Gröbner basis of $J_{G}$ with respect to a monomial order $<$ and set $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Then
$\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$
is the reduced Gröbner basis of $K_{G}$ with respect to $<$.
Proof.
Since $\mathcal{G}_{1}$ is a Gröbner basis of $J_{G}$, we have $J_{G}=\langle\mathcal{G}_{1}\rangle$. It follows from $M_{G}=\langle\mathcal{G}_{2}\rangle$ that ${\mathcal{G}}$ is a set of generators of $K_{G}=J_{G}+M_{G}$. We apply Buchberger Criterion to ${\mathcal{G}}$. If $p,q\in\mathcal{G}_{1}\setminus M_{G}$, then the remainder of $S(p,q)$ on division by ${\mathcal{G}}$ is 00 since $\mathcal{G}_{1}$ is a Gröbner basis of $J_{G}$. If $p,q\in\mathcal{G}_{2}$, then both $p$ and $q$ are monomials, and hence $S(p,q)=0$. Suppose that $p\in\mathcal{G}_{2}$ and $q\in\mathcal{G}_{1}\setminus M_{G}$. Let $p=x_{S}x_{T}$ with $S,T\in S(G)$ and $S\cap T\neq\emptyset$, and let $q={\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ where $f$ and $g$ are $k$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and ${\rm in}_{<}(q)={\mathbf{x}}_{f}$.
Case 1. ($p$ and ${\mathbf{x}}_{f}$ are relatively prime.) Then $S(p,q)=p{\mathbf{x}}_{g}$ is divided by $p\in\mathcal{G}_{2}$.
Case 2. (${\mathbf{x}}_{f}$ is divided by $x_{S}$.) From Lemma 6.3, ${\mathbf{x}}_{f}$ is not divided by $x_{T}$ since $q\notin M_{G}$. Then $S(p,q)=x_{T}{\mathbf{x}}_{g}$. Let $i\in S\cap T$. Since ${\mathbf{x}}_{f}$ is divided by $x_{S}$, $\pi({\mathbf{x}}_{f})=\pi({\mathbf{x}}_{g})$ is divided by $t_{i}$. Thus $\pi(x_{T}{\mathbf{x}}_{g})$ is divided by $t_{i}^{2}$. Hence $x_{T}{\mathbf{x}}_{g}$ is divided by a monomial $x_{S^{\prime}}x_{T}$ such that $i\in S^{\prime}\cap T$. Then the remainder of $S(p,q)=x_{T}{\mathbf{x}}_{g}$ on division by ${\mathcal{G}}_{2}$ is 00.
Therefore the remainder of any $S$-polynomial on division by ${\mathcal{G}}$ is 00. By Buchberger Criterion, ${\mathcal{G}}$ is a Gröbner basis of $K_{G}$. Since ${\mathcal{G}}_{1}$ is reduced, it is easy to see that ${\mathcal{G}}$ is reduced. ∎
Similarly to $\langle[I_{G}]_{2}\rangle$ and $J_{G}$, we can determine if two $k$-colorings of $G$ are Kempe equivalent by $K_{G}$.
Proposition 6.5.
Let $G$ be a graph on $[d]$ and let $f$ and $g$ be $k$-colorings of an induced subgraph of $G$. Then $f\sim_{k}g$ if and only if ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$.
Proof.
Let $f$ and $g$ be $k$-colorings of an induced subgraph $G_{0}$ of $G$. From Theorem 5.1 it is enough to show that ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$ if and only if ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$. Since $J_{G}\subset K_{G}$, ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$ if ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$.
Suppose that ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$. We show that ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$. Let $\mathcal{G}$ be the reduced Gröbner basis of $K_{G}$ with respect to a monomial order $<$. From Proposition 6.4, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ where $\mathcal{G}_{1}$ is the reduced Gröbner basis of $J_{G}$ with respect to $<$ and $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Let ${\mathbf{x}}_{f^{\prime}}$ (resp. ${\mathbf{x}}_{g^{\prime}}$) denote the remainder of ${\mathbf{x}}_{f}$ (resp. ${\mathbf{x}}_{g}$) on division by $\mathcal{G}_{1}\setminus M_{G}$. Suppose that ${\mathbf{x}}_{f^{\prime}}\neq{\mathbf{x}}_{g^{\prime}}$. Since $f$ and $g$ are $k$-colorings of $G_{0}$, we have ${\mathbf{x}}_{f},{\mathbf{x}}_{g}\notin M_{G}$ from Lemma 6.1. From Lemma 6.3, ${\mathbf{x}}_{f^{\prime}},{\mathbf{x}}_{g^{\prime}}\notin M_{G}$ since both ${\mathbf{x}}_{f}-{\mathbf{x}}_{f^{\prime}}$ and ${\mathbf{x}}_{g}-{\mathbf{x}}_{g^{\prime}}$ belong to $J_{G}$. Thus ${\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{g^{\prime}}$ ($\neq 0$) is the normal form of ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$ with respect to $\mathcal{G}$. This contradicts that $\mathcal{G}$ is a Gröbner basis of $K_{G}$. Thus ${\mathbf{x}}_{f^{\prime}}={\mathbf{x}}_{g^{\prime}}$. Then the normal form of ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ with respect to $\mathcal{G}_{1}\setminus M_{G}$ is zero, and hence ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$. ∎
Now, we give a proof of Theorem 1.2. In fact, Theorem 1.2 follows from the following.
Theorem 6.6.
Let $G$ be a graph on $[d]$ and $<$ a monomial order on $R[G]$, and let $\{{\mathbf{x}}_{f_{1}},\dots,{\mathbf{x}}_{f_{s}}\}$ be the set of all standard monomials of degree $k$ with respect to ${\rm in}_{<}(K_{G})$. Then each $f_{i}$ is a $k$-coloring of an induced subgraph of $G$. In addition, given an induced subgraph $G^{\prime}$ of $G$,
$\{f_{1},\dots,f_{s}\}\cap\mathcal{C}_{k}(G^{\prime})$
is a complete representative system for ${\rm kc}(G^{\prime},k)$.
Remark 6.7.
If ${\mathbf{x}}_{f}=x_{S_{1}}x_{S_{2}}\cdots x_{S_{k}}$ is a standard monomial of degree $k$ with respect to ${\rm in}_{<}(K_{G})$, then $f$ is a $k$-coloring of the induced subgraph of $G$ on $\bigcup_{i=1}^{k}S_{i}$.
Proof of Theorem 6.6.
Let $\mathcal{G}$ be the reduced Gröbner basis of $K_{G}$ with respect to a monomial order $<$. From Proposition 6.4, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ where $\mathcal{G}_{1}$ is the reduced Gröbner basis of $J_{G}$ with respect to $<$ and $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Since each ${\mathbf{x}}_{f_{i}}$ is not divided by any monomial in $\mathcal{G}_{2}$, from Lemma 6.1, each $f_{i}$ is a $k$-coloring of an induced subgraph of $G$.
Suppose that $f_{i}\sim_{k}f_{j}$ for some $1\leq i<j\leq s$. From Theorem 5.1, ${\mathbf{x}}_{f_{i}}-{\mathbf{x}}_{f_{j}}$ belongs to $J_{G}(\subset K_{G})$. Then ${\rm in}_{<}({\mathbf{x}}_{f_{i}}-{\mathbf{x}}_{f_{j}})$ is not standard, a contradiction. Hence $f_{i}\not\sim_{k}f_{j}$ for any $1\leq i<j\leq s$.
Let $f$ be a $k$-coloring of $G^{\prime}$, and let ${\mathbf{x}}_{f^{\prime}}$ be the remainder of ${\mathbf{x}}_{f}$ on division by $\mathcal{G}_{1}\setminus M_{G}$. Then ${\mathbf{x}}_{f}-{\mathbf{x}}_{f^{\prime}}$ belongs to $J_{G}$. Since ${\mathbf{x}}_{f}\notin M_{G}$, we have ${\mathbf{x}}_{f^{\prime}}\notin M_{G}$ from Lemma 6.3. Thus ${\mathbf{x}}_{f^{\prime}}$ is the normal form of ${\mathbf{x}}_{f}$ with respect to $\mathcal{G}$, and hence ${\mathbf{x}}_{f^{\prime}}={\mathbf{x}}_{f_{j}}$ for some $j$. Since ${\mathbf{x}}_{f}-{\mathbf{x}}_{f_{j}}$ belongs to $J_{G}$, we have $f\sim_{k}f_{j}$ from Theorem 5.1, ∎
We see an example of Theorem 6.6.
Example 6.8.
Let $G$ be the graph as in Example 5.3. We consider the reverse lexicographic order $<$ on $R[G]$ such that
$x_{\emptyset}<x_{\{1\}}<\cdots<x_{\{6\}}<x_{\{1,5\}}<x_{\{1,6\}}<x_{\{2,4\}}<x_{\{2,6\}}<x_{\{3,4\}}<x_{\{3,5\}}.$
Then there are $65$ standard monomials of degree $3$ with respect to ${\rm in}_{<}(K_{G})$. In particular, the standard monomials $x_{\{1,5\}}x_{\{2,6\}}x_{\{3,4\}}$ and $x_{\{1,6\}}x_{\{2,4\}}x_{\{3,5\}}$ correspond to $3$-colorings of $G$. It then follows from Theorem 6.6 that the associated $3$-colorings, which are the $3$-colorings as in Example 5.3, form a complete representative system for ${\rm kc}(G,k)$.
As a consequence of Theorem 6.6, the number of $k$-Kempe classes ${\rm Kc}(G,k)$ can be computed by using Hilbert functions. We denote ${\rm Ind}(G)$ the set of all induced subgraphs of $G$ and denote ${\rm Ind}_{m}(G)$ the set of all induced subgraphs of $G$ with $m$ vertices.
Corollary 6.9.
Let $G$ be a graph on $[d]$. Then one has
$H(R[G]/K_{G},k)=\sum_{G^{\prime}\in{\rm Ind}(G)}{\rm Kc}(G^{\prime},k).$
In particular,
${\rm Kc}(G,k)=\sum_{m=0}^{d}(-1)^{d-m}\sum_{G^{\prime}\in{\rm Ind}_{m}(G)}H(R[G^{\prime}]/K_{G^{\prime}},k).$
Example 6.10.
Let $G$ be the graph as in Example 5.3. Then one has
$H(R[G]/K_{G},k)=\left\{\begin{array}[]{cc}1&k=0,\\ 13&k=1,\\ 49&k=2,\\ 65&k=3,\\ 64&k\geq 4,\end{array}\right.$
and
$H(R[G^{\prime}]/K_{G^{\prime}},3)=32$
for any $G^{\prime}\in{\rm Ind}_{5}(G)$. Note that ${\rm Kc}(G^{\prime},k)\geq 1$ for any $k\geq\chi(G)$ and $G^{\prime}\in{\rm Ind}(G)$. Moreover, one has $|{\rm Ind}(G)|=2^{6}=64$. Hence we have the following from $H(R[G]/K_{G},k)$.
  • Since $H(R[G]/K_{G},2)<2^{6}$, one has $\chi(G)\geq 3$. In fact, $\chi(G)=3$ in this case.
  • Let $k\geq 4$. Since $H(R[G]/K_{G},k)=2^{6}$, ${\rm Kc}(G^{\prime},k)=1$ for any $G^{\prime}\in{\rm Ind}(G)$.
  • Since $H(R[G^{\prime}]/K_{G^{\prime}},3)=2^{5}$, ${\rm Kc}(G^{\prime\prime},3)=1$ for any $G^{\prime\prime}\in{\rm Ind}(G)$ with $G\neq G^{\prime\prime}$. Hence one has ${\rm Kc}(G,3)=2$ from $H(R[G]/K_{G},3)=65$.
Let $I$ be a graded ideal of $R$. In general, the Hilbert function $H(R/I,k)$ is not always a polynomial. However, there exists a unique polynomial $P_{R/I}\in{\mathbb{Q}}[k]$ such that $H(R/I,k)=P_{R/I}(k)$ for $k$ large enough. We call $P_{R/I}$ the Hilbert polynomial of $R/I$. We show that $P_{R[G]/K_{G}}$ is a constant which depends only on the number of vertices.
Proposition 6.11.
Let $G$ be a graph on $[d]$. Then one has
$P_{R[G]/K_{G}}(k)=2^{d}.$
Proof.
Let $\Delta$ be the maximum degree of $G$. Then it follows from [5, Corollary 2.5] that for any integer $k\geq\Delta+1$ and for any induced subgraph $G^{\prime}$ of $G$, one has ${\rm Kc}(G^{\prime},k)=1$. Hence for any integer $k\geq\Delta+1$, we obtain
$H(R[G]/K_{G},k)=\sum_{G^{\prime}\in{\rm Ind}(G)}{\rm Kc}(G^{\prime},k)=\sum_{G^{\prime}\in{\rm Ind}(G)}1=|{\rm Ind}(G)|=2^{d}.$
This implies that
$P_{R[G]/K_{G}}(k)=2^{d},$
as desired. ∎
7. Algorithms
In this section, by using Theorems 1.1 and 1.2 and techniques on Gröbner bases, we introduce several algorithms related to Kempe equivalence.
First we give Algorithm 7.1 to compute the reduced Gröbner basis of $K_{G}$. From Proposition 4.2, ${\mathcal{G}}_{2}$ in Algorithm 7.1 is a Gröbner basis of $I_{G}$. It then follows that ${\mathcal{G}}_{3}$ in Algorithm 7.1 is a system of generators of $J_{G}$.
Algorithm 7.1 Computation of the reduced Gröbner basis of $K_{G}$
0:  The set of all stable sets of $G$, and a monomial order $<$ on $R[G]$.
0:  The reduced Gröbner basis of $K_{G}$ with respect to $<$.
1:  :=${\mathcal{F}}:=\{x_{S\setminus\{i\}}x_{\{i\}}-x_{S}x_{\emptyset}\mid i\in S\in S(G),\ |S|\geq 2\}$.
2:  From ${\mathcal{F}}$, compute the reduced Gröbner basis ${\mathcal{G}}_{1}$ of $L_{G}$ with respect to a reverse lexicographic order such that $x_{S}\geq x_{\emptyset}$ for any $S\in S(G)$.
3:  $\mathcal{G}_{2}:=\{g/x_{\emptyset}^{k}\mid g\in\mathcal{G},k\in{\mathbb{Z}}_{\geq 0}$, $x_{\emptyset}^{k}$ divides $g$, $x_{\emptyset}^{k+1}$ does not divide $g\}$.
4:  $\mathcal{G}_{3}:=\{x_{S_{1}}x_{S_{2}}-x_{S_{3}}x_{S_{4}}\in{\mathcal{G}}_{2}\mid S_{1},S_{2},S_{3},S_{4}\in S(G),\ S_{1}\cap S_{2}=S_{3}\cap S_{4}=\emptyset\}$.
5:  Compute the reduced Gröbner basis ${\mathcal{G}}$ of $K_{G}$ with respect to $<$ from a system of generators ${\mathcal{G}}_{3}\cup\{x_{S}x_{T}\mid S,T\in S(G),S\cap T\neq\emptyset\}$.
6:  return  ${\mathcal{G}}$.
Using the reduced Gröbner basis of $K_{G}$, Algorithm 7.2 determines whether $f\sim_{k}g$ or not. The correctness of Algorithm 7.2 is an immediate consequence of Lemma 3.3 and Proposition 6.5. From Theorem 5.1, it is possible to replace $K_{G}$ with either $\langle[I_{G}]_{2}\rangle$ or $J_{G}$ in Algorithm 7.2.
Algorithm 7.2 Determination of the Kempe equivalence
0:  $k$-colorings $f$ and $g$ of $G$, and the reduced Gröbner basis $\mathcal{G}$ of $K_{G}$.
0:  “$f\sim_{k}g$” or “$f\not\sim_{k}g$”.
1:  Compute the normal form $m$ of ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ with respect to $\mathcal{G}$.
2:  if $m=0$ then
3:     return  “$f\sim_{k}g$
4:  else
5:     return  “$f\not\sim_{k}g$
6:  end if
On the other hand, by the reduced Gröbner basis of $K_{G}$, we can compute the set of all standard monomials of degree $k$ with respect to the initial ideal ${\rm in}_{<}(K_{G})$. From Theorem 1.2, we have Algorithm 7.3 to compute a complete representative system for ${\rm kc}(G,k)$.
Algorithm 7.3 Computation of a complete representative system for ${\rm kc}(G,k)$
0:  The reduced Gröbner basis $\mathcal{G}$ of $K_{G}$.
0:  A complete representative system for ${\rm kc}(G,k)$.
1:  From ${\mathcal{G}}$, compute the set $\{{\mathbf{x}}_{f_{1}},\dots,{\mathbf{x}}_{f_{s}}\}$ of all standard monomials of degree $k$ with respect to the initial ideal ${\rm in}_{<}(K_{G})$.
2:  $C:=\{\}.$
3:  for $i=1,2,\dots,s$ do
4:     if ${\mathbf{x}}_{f_{i}}=x_{S_{1}}\cdots x_{S_{k}}$ satisfies $[d]=S_{1}\cup\cdots\cup S_{k}$ then
5:        $C:=C\cup\{f_{i}\}$.
6:     end if
7:  end for
8:  return  $C$
Algorithm 7.4 enumerates all elements in a Kempe equivalent class. The correctness of Algorithm 7.4 is guaranteed by Theorem 6.6 together with the fact that, with respect to a Gröbner basis, the normal form of the monomials in the same residue class is unique. In addition, Algorithm 7.4 is based on [8, Algorithm 5.7] for enumeration of fibers using a Gröbner basis of a toric ideal. As stated in [8], “reverse search” technique is useful to improve the efficiency.
Algorithm 7.4 Enumeration of elements in a Kempe equivalent class
0:  $k$-coloring $f$ of $G$, and the reduced Gröbner basis $\mathcal{G}$ of $K_{G}$.
0:  All $k$-colorings of $G$ which are Kempe equivalent to $f$ (up to permutations of colors).
1:  Compute the normal form ${\mathbf{x}}_{f^{\prime}}$ of ${\mathbf{x}}_{f}$ with respect to $\mathcal{G}$.
2:  $A:=\{f^{\prime}\}$, $B:=\{\}$.
3:  while $A\neq\{\}$ do
4:     Choose $g\in A$.
5:     for ${\mathbf{x}}_{g_{1}}-{\mathbf{x}}_{g_{2}}\in{\mathcal{G}}$ with ${\mathbf{x}}_{g_{1}}>{\mathbf{x}}_{g_{2}}$ do
6:        if ${\mathbf{x}}_{g_{2}}$ divides ${\mathbf{x}}_{g}$ and $g^{\prime}\notin B$ where ${\mathbf{x}}_{g^{\prime}}={\mathbf{x}}_{g}{\mathbf{x}}_{g_{1}}/{\mathbf{x}}_{g_{2}}$ then
7:           $A:=A\cup\{g^{\prime}\}$.
8:        end if
9:     end for
10:     $A:=A\setminus\{g\}$, $B:=B\cup\{g\}$.
11:  end while
12:  return  $B$.
Finally, we give an algorithm to find a sequence of Kempe switchings. We define a Kempe basis which is the set of sequences of colorings corresponding to the reduced Gröbner basis of $K_{G}$.
Definition 7.1.
Work with the same notation as in Theorem 6.4, that is, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ is the reduced Gröbner basis of $K_{G}$ with respect to $<$. Let $\mathcal{G}_{1}\setminus M_{G}=\{{\mathbf{x}}_{p_{1}}-{\mathbf{x}}_{q_{1}},\dots,{\mathbf{x}}_{p_{t}}-{\mathbf{x}}_{q_{t}}\}$. From Theorem 1.1, $p_{j}$ and $q_{j}$ are Kempe equivalent for each $j$. Then
$\left\{(f_{1}^{(1)},\dots,f_{s_{1}}^{(1)}),\dots,(f_{1}^{(t)},\dots,f_{s_{t}}^{(t)})\right\}$
is called a Kempe basis of $G$ with respect to $<$ if $f_{1}^{(j)}=p_{j}$, $f_{s_{j}}^{(j)}=q_{j}$, and $f_{i}^{(j)}$ is obtained from $f_{i-1}^{(j)}$ by a Kempe switching.
In order to give an algorithm to find a Kempe basis, the following three Procedures are important.
Procedure 1. (${\mathbf{x}}_{f}-{\mathbf{x}}_{g}=x_{S_{1}}x_{S_{2}}-x_{S_{3}}x_{S_{4}}\in J_{G}\mapsto$ a sequence of Kempe switching from $f$ to $g$.)
Suppose that $f$ and $g$ are 2-colorings of an induced subgraph $G_{0}$ of $G$. Let ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}=x_{S_{1}}x_{S_{2}}-x_{S_{3}}x_{S_{4}}\in J_{G}$. Then $S_{1}\cap S_{2}=S_{3}\cap S_{4}=\emptyset$ and $S_{1}\cup S_{2}=S_{3}\cup S_{4}$. Hence $g$ is obtained from $f$ by setting
$g(x)=\begin{cases}f(x)&x\notin H,\\ 1&x\in H\mbox{ and }f(x)=2,\\ 2&x\in H\mbox{ and }f(x)=1,\end{cases}$
where $H$ is the induced subgraph of $G_{0}$ on the vertex set $(S_{1}\setminus S_{3})\sqcup(S_{2}\setminus S_{4})$. Suppose that $H$ has $p$ connected components $H_{1},\dots,H_{p}$. For $l=1,2,\dots,p$, let $f_{0}=f$ and let
$f_{l}(x)=\begin{cases}f(x)&x\notin H,\\ 1&x\in H_{1}\cup\dots\cup H_{l}\mbox{ and }f(x)=2,\\ 2&x\in H_{1}\cup\dots\cup H_{l}\mbox{ and }f(x)=1.\end{cases}$
Then $f_{0},f_{1},\ldots,f_{p}$ is a sequence of $2$-colorings of $G_{0}$ such that $f_{0}=f$, $f_{p}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching.
Procedure 2. Let $f^{\prime}$ and $g^{\prime}$ be $k$-colorings of an induced subgraph $G_{1}$ of $G$. Suppose that ${\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{g^{\prime}}={\bf x}_{h}({\mathbf{x}}_{f}-{\mathbf{x}}_{g})$. Then $f$ and $g$ are the restriction of $f^{\prime}$ and $g^{\prime}$ to some induced subgraph $G_{2}$ of $G_{1}$, respectively. Hence if $f_{0},f_{1},\ldots,f_{p}$ is a sequence of $k^{\prime}$-colorings of $G_{2}$ such that $f_{0}=f$, $f_{p}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching, then $f_{0}^{\prime},f_{1}^{\prime},\ldots,f_{p}^{\prime}$ satisfy $f_{0}^{\prime}=f^{\prime}$, $f_{p}^{\prime}=g^{\prime}$, and $f_{i}^{\prime}$ is obtained from $f_{i-1}^{\prime}$ by a Kempe switching, where $f_{i}^{\prime}$ is a $k$-coloring of $G_{1}$ obtained by combining $f_{i}$ with $h$.
Procedure 3. Let ${\mathbf{x}}_{p}-{\mathbf{x}}_{q}$ be a binomial in the reduced Gröbner basis ${\mathcal{G}}$ of $K_{G}$ with respect to a monomial order $<$. From Proposition 6.4, any binomial ${\mathbf{x}}_{p}-{\mathbf{x}}_{q}\in{\mathcal{G}}$ is obtained from the reduced Gröbner basis of $J_{G}$ with respect to $<$ which is computed from
$\{{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are $2$-colorings of an induced subgraph of $G$}\}$
by Buchberger Algorithm. By keeping track of the computation of Buchberger Algorithm, we can compute an expression
${\mathbf{x}}_{p}-{\mathbf{x}}_{q}=\sum_{r=1}^{s}{\bf x}_{w_{r}}({\mathbf{x}}_{f_{r}}-{\mathbf{x}}_{g_{r}}),$
where $f_{r}$ and $g_{r}$ are $2$-colorings of an induced subgraph of $G$ for each $r$,
${\bf x}_{w_{r}}{\mathbf{x}}_{g_{r}}={\bf x}_{w_{r+1}}{\mathbf{x}}_{f_{r+1}}$
for each $r=1,\dots,s-1$, and ${\mathbf{x}}_{p}={\bf x}_{w_{1}}{\mathbf{x}}_{f_{1}}$, ${\mathbf{x}}_{q}={\bf x}_{w_{s}}{\mathbf{x}}_{g_{s}}$. See [1, Chapter 2, §1] for details. Then we can compute a sequence of Kempe switchings from $u_{r}$ to $v_{r}$ where ${\mathbf{x}}_{u_{r}}={\mathbf{x}}_{w_{r}}{\mathbf{x}}_{f_{r}}$ and ${\mathbf{x}}_{v_{r}}={\mathbf{x}}_{w_{r}}{\mathbf{x}}_{g_{r}}$ by Procedures 1 and 2. Combining them, we have a sequence of Kempe switchings from $p$ to $q$.
Using Procedures 1, 2, and 3, we have Algorithm 7.5 that computes a Kempe basis. Note that, if the reduced Gröbner basis of $K_{G}$ consists of quadratic polynomials, then we can skip large part of the procedure in Algorithm 7.5.
Algorithm 7.5 Construction of a Kempe basis
0:  A monomial order $<$ on $R[G]$.
0:  A Kempe basis of $G$ with respect to $<$.
1:  Compute a system of (quadratic) generators
=${\mathcal{F}}=\{{\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g_{1}},\dots,{\mathbf{x}}_{f_{t}}-{\mathbf{x}}_{g_{t}}\}\cup\{x_{S}x_{T}\mid S,T\in S(G),S\cap T\neq\emptyset\}$
of $K_{G}$.
2:  for $r=1,2,\dots,t$ do
3:     By Procedure 1, compute a sequence $F_{r}$ of 2-colorings corresponding to a sequence of Kempe switchings from $f_{r}$ to $g_{r}$.
4:  end for
5:  Compute the reduced Gröbner basis ${\mathcal{G}}$ of $K_{G}$ with respect to $<$ from ${\mathcal{F}}$.
6:  $A:=\{\}$.
7:  for ${\mathbf{x}}_{p}-{\mathbf{x}}_{q}\in\mathcal{G}$ do
8:     By keeping track of the computation of $\mathcal{G}$ from ${\mathcal{F}}$, compute the expression
${\mathbf{x}}_{p}-{\mathbf{x}}_{q}=\sum_{r=1}^{s}{\bf x}_{w_{r}}({\mathbf{x}}_{f_{i_{r}}}-{\mathbf{x}}_{g_{i_{r}}}),$
where ${\bf x}_{w_{r}}{\mathbf{x}}_{g_{i_{r}}}={\bf x}_{w_{r+1}}{\mathbf{x}}_{f_{i_{r+1}}}$ for each $r=1,\dots,s-1$, and ${\mathbf{x}}_{p}={\bf x}_{w_{1}}{\mathbf{x}}_{f_{i_{1}}}$, ${\mathbf{x}}_{q}={\bf x}_{w_{s}}{\mathbf{x}}_{g_{i_{s}}}$.
9:     for $r=1,2,\dots,s$ do
10:        $f:=f_{i_{r}}^{\prime}$ where ${\mathbf{x}}_{f_{i_{r}}^{\prime}}={\bf x}_{w_{r}}{\mathbf{x}}_{f_{i_{r}}}$, and $g:=g_{i_{r}}^{\prime}$ where ${\mathbf{x}}_{g_{i_{r}}^{\prime}}={\bf x}_{w_{r}}{\mathbf{x}}_{g_{i_{r}}}$.
11:        By extending $F_{i_{r}}$, find a sequence of colorings $h_{1}^{(r)},\dots,h_{t_{r}}^{(r)}$ where $h_{1}^{(r)}=f$, $h_{t_{r}}^{(r)}=g$ and $h_{j}^{(r)}$ is obtained from $h_{j-1}^{(r)}$ by a Kempe switching.
12:     end for
13:     $A:=A\cup\{(h_{1}^{(1)},\dots,h_{t_{1}}^{(1)}(=h_{1}^{(2)}),h_{2}^{(2)},\dots,h_{t_{2}}^{(2)}(=h_{1}^{(3)}),\dots,h_{t_{s}}^{(s)})\}$
14:  end for
15:  return  $A$
We now use a Kempe basis to find a sequence of Kempe switchings between two colorings. Suppose that the normal form ${\mathbf{x}}_{f^{\prime}}$ of ${\mathbf{x}}_{f}$ with respect to the reduced Gröbner basis $\mathcal{G}$ of $K_{G}$ is given by
${\mathbf{x}}_{f}={\mathbf{x}}_{f_{0}}\underset{h_{1}}{\rightarrow}{\mathbf{x}}_{f_{1}}\underset{h_{2}}{\rightarrow}\cdots\underset{h_{t}}{\rightarrow}{\mathbf{x}}_{f_{t}}={\mathbf{x}}_{f^{\prime}},\ \ \ (h_{j}\in\mathcal{G}).$
By Procedure 2, we can construct a sequence of Kempe switchings from $f_{i-1}$ to $f_{i}$ by extending a sequence of Kempe switchings corresponding to $h_{i-1}$ in a Kempe basis. Using this fact, we have Algorithm 7.6 to find a sequence of Kempe switchings.
Algorithm 7.6 Construction of a sequence of Kempe switchings
0:  $k$-colorings $f$ and $g$ of $G$, the reduced Gröbner basis $\mathcal{G}$ of $K_{G}$, and a Kempe basis $\mathcal{K}$ of $G$.
0:  “a sequence of Kempe switchings from $f$ to $g$” or “$f\not\sim_{k}g$”.
1:  Compute the normal form ${\mathbf{x}}_{f^{\prime}}$ of ${\mathbf{x}}_{f}$ with respect to $\mathcal{G}$.
2:  Compute the normal form ${\mathbf{x}}_{g^{\prime}}$ of ${\mathbf{x}}_{g}$ with respect to $\mathcal{G}$.
3:  if ${\mathbf{x}}_{f^{\prime}}={\mathbf{x}}_{g^{\prime}}$ then
4:     By Procedure 2, from $\mathcal{K}$, find a sequence of colorings $f=f_{1},\dots,f_{s},f^{\prime}$, where $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching.
5:     By Procedure 2, from $\mathcal{K}$, find a sequence of colorings $g=g_{1},\dots,g_{t},g^{\prime}$ ($=f^{\prime}$), where $g_{i}$ is obtained from $g_{i-1}$ by a Kempe switching.
6:     return  $f_{1},\dots,f_{s},f^{\prime},g_{t},g_{t-1},\dots,g_{1}$.
7:  else
8:     return  “$f\not\sim_{k}g$
9:  end if
Acknowledgment
The second author was partially supported by JSPS KAKENHI 22K13890.